OpenStudy (anonymous):

find the intercepts of -2y+x=2y-8

OpenStudy (solomonzelman):

\[-2y+x=2y-8\]\[-2y+x\color{red} {-2y-x}=2y-8\color{red} {-2y-x}\]\[-4y=-8-x\]\[-4y\color{red} {/(-4)}=-x\color{red} {/(-4)}-8\color{red} {/(-4)}\]\[y=\frac{1}{4}x+2\]

OpenStudy (solomonzelman):

Y-INTERCEPT. \[y=\frac{1}{4}\color{red} {(0)}+2~~~~~~->~~~~~~y=2\] X-INTERCEPT. \[\color{red} {(0)}=\frac{1}{4}x+2\]\[-\frac{1}{4}x=2\]\[-\frac{1}{4}x\color{red} {\times (-4)}=2\color{red} {\times (-4)}\]\[x=-8\]